# LUSIFER: Language Universal Space Integration for Enhanced Multilingual Embeddings with Large Language Models

### v1

@@ Abstract
Recent advancements in large language models (LLMs) based embedding models have established new state-of-the-art benchmarks for text embedding tasks, particularly in dense vector-based retrieval. However, these models predominantly focus on English, leaving multilingual embedding capabilities largely unexplored. To address this limitation, we present LUSIFER, a novel zero-shot approach that adapts LLM-based embedding models for multilingual tasks without requiring multilingual supervision. LUSIFER’s architecture combines a multilingual encoder, serving as a language-universal learner, with an LLM-based embedding model optimized for embedding-specific tasks. These components are seamlessly integrated through a minimal set of trainable parameters that act as a connector, effectively transferring the multilingual encoder’s language understanding capabilities to the specialized embedding model. Additionally, to comprehensively evaluate multilingual embedding performance, we introduce a new benchmark encompassing 5 primary embedding tasks, 123 diverse datasets, and coverage across 14 languages. Extensive experimental results demonstrate that LUSIFER significantly enhances the multilingual performance across various embedding tasks, particularly for medium and low-resource languages, without requiring explicit multilingual training data.
LUSIFER: Language Universal Space Integration for Enhanced Multilingual Embeddings with Large Language Models
Hieu Man1, Nghia Trung Ngo1, Viet Dac Lai2, Ryan A. Rossi2, Franck Dernoncourt2, Thien Huu Nguyen1 1Dept. of Computer Science, University of Oregon, OR, USA 2Adobe Research, USA {hieum,nghian,thienn}@uoregon.edu {viet.lai,ryrossi,franck.dernoncourt}@adobe.com
1 Introduction
@@
Text embeddings, which provide dense vector representations of textual content Mikolov et al. (2013); Devlin et al. (2019), have become fundamental building blocks in modern natural language processing. These embeddings encode semantic information and serve as an important component for numerous downstream applications, ranging from information retrieval and document reranking to classification, clustering, and semantic textual similarity assessment. Recently, the significance of high-quality embeddings has been further amplified by their crucial role in retrieval-augmented generation (RAG) systems Lewis et al. (2020b). RAG architectures enable large language models (LLMs) to dynamically access and integrate external or proprietary knowledge without the need for model parameter updates, substantially enhancing their adaptability and accuracy Wang et al. (2023); Liu et al. (2024b); Gao et al. (2024).
@@
The evolution of embedding models has witnessed remarkable advancements, progressing from static word embeddings Robertson et al. (2009) through contextualized representations Reimers and Gurevych (2019); Gao et al. (2021b); Ni et al. (2021a) to state-of-the-art LLM-based embedding models Wang et al. (2024b) that harness the sophisticated semantic understanding capabilities of large language models. These developments have substantially enhanced performance across various embedding tasks Luo et al. (2024), achieving unprecedented accuracy in semantic similarity and retrieval applications. However, a critical limitation remains: the predominant focus on English in LLM-based embedding models has created a significant disparity in multilingual capabilities. This gap is especially pronounced in medium and low-resource languages, where English-centric models exhibit substantial performance degradation due to insufficient language-specific training data Wang et al. (2020); Thakur et al. (2024). While recent advances in multilingual embedding models, particularly those leveraging multilingual pre-trained architectures, have demonstrated promising results in multilingual embedding tasks Li et al. (2023); Wang et al. (2024c); Chen et al. (2024), their reliance on explicit multilingual supervision for embeddings constrains their applicability primarily to languages with abundant training resources, leaving the challenge of true language-agnostic representation largely unaddressed.
To address this challenge, we present LUSIFER, a novel zero-shot approach that adapts English LLM-based embedding models for multilingual tasks without requiring explicit multilingual supervision. Drawing inspiration from recent advances in multimodal integration Liu et al. (2024a); Lu et al. (2024), LUSIFER employs a unique architecture that bridges the gap between multilingual understanding and specialized embedding capabilities. At its core, LUSIFER leverages the robust multilingual representations from XLM-R Conneau et al. (2020) and introduces a learnable connector mechanism to interface with English-optimized LLM embedding models. This approach enables LUSIFER to effectively transfer the multilingual understanding of XLM-R to the target LLM while inheriting advanced embedding capabilities of the LLM. In this way, LUCIFER can achieve effective multilingual representation capabilities without requiring explicit multilingual training data.
@@
We conduct comprehensive evaluations of LUSIFER through extensive experiments across 123 diverse datasets spanning 14 languages, focusing on five fundamental embedding tasks: Classification, Clustering, Reranking, Retrieval, and Semantic Textual Similarity (STS). Our experimental results demonstrate that LUSIFER substantially enhances the performance of English-centric LLM-based embedding models, achieving average improvements of 3.19 points across all tasks, with particularly significant gains observed for medium and low-resource languages (up to 22.15 improvement). To validate LUSIFER’s broader applicability and cross-lingual capabilities, we extend our evaluation to cross-lingual tasks using four comprehensive datasets that encompass over 100 languages, including several critically low-resource languages. LUSIFER significantly outperforms existing English-centric embedding models by 5.75 on average in cross-lingual scenarios. These results demonstrate the effectiveness of our approach in enhancing multilingual representation capabilities without explicit multilingual supervision.
@@
The theoretical foundation for LUSIFER’s effectiveness lies in its ability to create a language-agnostic universal space through the integration of a multilingual encoder Pires et al. (2019); Libovický et al. (2020). We hypothesize that this universal space serves as a bridge between different languages, enabling the target language model to process semantic information independently of the input language. By mapping these language-neutral representations to the target model’s input space, we conjecture that the target LLM can grasp the semantics of these representations, thereby improving the quality of output embeddings across multiple languages. This mechanism allows the model to become less dependent on the specific language of the input, enabling it to better capture semantic information for embedding tasks in languages it rarely encountered during pretraining. Our empirical analysis using t-SNE visualization supports this hypothesis.
2 Related Work
2.1 English-centric Embedding Models
Text embedding models have experienced significant advancement in recent years, driven by the evolution of pre-trained language models. Early successes with BERT-based architectures, as demonstrated in Sentence-BERT Reimers and Gurevych (2019), SimCSE Gao et al. (2021b), and DPR Karpukhin et al. (2020), established the foundation for modern embedding approaches. The field has since progressed to leverage LLMs, with recent works Wang et al. (2024b); Muennighoff et al. (2024); BehnamGhader et al. (2024); Lee et al. (2024); Man et al. (2024) demonstrating substantial improvements in embedding quality and task performance through the enhanced representational capacity of LLMs Luo et al. (2024). However, these advances primarily benefit high-resource language applications, as most state-of-the-art LLM-based embedding models are derived from English-centric foundation models Jiang et al. (2023); Touvron et al. (2023) and trained predominantly on English or high-resource language datasets Wang et al. (2024b). This bias has resulted in a significant performance gap between high-resource and low-resource languages, limiting the global applicability of these models. Our proposed method, LUSIFER, addresses this limitation by enabling effective multilingual representation without multilingual training data.
2.2 Zero-shot Multilingual Embedding
Multilingual Embedding has evolved through several distinct methodological approaches, each addressing the fundamental challenge of bridging language gaps in embedding tasks. Early successful approaches relied on translation models to enable multilingual understanding Liu et al. (2020); Shi et al. (2021); Zhang and Misra (2022). While effective, these methods introduced operational complexity by requiring external translation systems, limiting their practical deployment and scalability.
The emergence of multilingual pre-trained language models, particularly XLM-R Conneau et al. (2020), opened new possibilities for multilingual transfer. Recent works have demonstrated promising results by fine-tuning such models with contrastive learning objectives on multilingual data Wang et al. (2024c); Chen et al. (2024); Sturua et al. (2024). However, these approaches face two key limitations: they require substantial multilingual training data, and moreover, they do not exploit the sophisticated semantic representations afforded by contemporary English-centric LLM architectures, which have demonstrated superior performance in capturing nuanced semantic relationships.
Recent advances in aligning multilingual and English-centric representations could offer a solution. By combining independently pre-trained representations, a paradigm that has shown remarkable success in multimodal alignment research Alayrac et al. (2022); Liu et al. (2024a); Lu et al. (2024), these works bridge the gap between visual encoders and language models to enhance visual comprehension. As such, similar principles can be applied to align multilingual representations with LLM-based semantic spaces. While related efforts have explored aligning multiple LLMs for improved reasoning capabilities in multilingual settings Bansal et al. (2024); Yoon et al. (2024), these approaches primarily target generation tasks and typically require large-scale alignment data. Our work extends these efforts by focusing on embedding tasks and leveraging a minimal set of parameters to align multilingual and English-centric representations, enabling enhanced multilingual representation capabilities without requirement for large-scale multilingual training data.
2.3 Multilingual Embedding Benchmarks
The evaluation landscape for multilingual embedding models has historically been fragmented across various benchmarks, each with significant limitations. While existing benchmarks have made valuable contributions, they often exhibit constrained scope: MINERS Winata et al. (2024) provides evaluation across multiple languages but is limited to classification and STS tasks with only 11 datasets; XNLI Conneau et al. (2018), XQuAD Artetxe et al. (2020), and SIB-200 Adelani et al. (2024) offer broad language coverage but focus exclusively on classification tasks; and MTEB Muennighoff et al. (2023), despite its diverse task selection, primarily addresses high-resource languages. To address these limitations, we introduce a comprehensive evaluation framework that encompasses 5 fundamental embedding tasks—Classification, Clustering, Reranking, Retrieval, and STS—across an extensive collection of 123 datasets spanning 14 languages. This holistic approach enables systematic evaluation across both task and language dimensions, providing unprecedented insights into models’ multilingual capabilities. Furthermore, our benchmark extends beyond traditional multilingual evaluation by incorporating cross-lingual tasks, featuring coverage of over 100 languages, including critically low-resource languages that have been historically underrepresented in existing benchmarks. This extensive coverage allows for a more nuanced understanding of embedding models’ performance across the global linguistic landscape.
@@ 2.3 Multilingual Embedding Benchmarks
Figure 1: Overview of LUSIFER. Left: Align a multilingual encoder with the target English-centric LLM only using English data and a minimal set of trainable parameter. Center: End-to-end representation finetune through contrastive learning on English text-embedding tasks using LoRA. Right: During inference, LUSIFER successfully processes text-embedding tasks across multiple languages.
3 Methodology
@@
Previous works demonstrate that representations of multilingual encoder models exhibit inherent language-agnostic properties, facilitating zero-shot multilingual transfer Pires et al. (2019); Libovický et al. (2020). Building upon this foundation, we propose LUSIFER, an embedding framework that aligns a multilingual encoder model with a target English-centric LLM’s representational space, enabling the target to encode semantics across multiple languages without extensive multilingual training. This section details our architectural design and two-stage training process for LUSIFER.
3.1 Model Architecture
@@
Following successful approaches in multimodal alignment Alayrac et al. (2022); Liu et al. (2024a); Lu et al. (2024), we implement the connector as a 2-layers feed-forward network, $\mathbf{FF}$, augmented with a single trainable token appended to the multilingual encoder’s hidden states. Formally, given input tokens $\mathbf{X}_{input}$ (with necessary padding), the multilingual encoder’s hidden states $\mathbf{H}_{enc}$ are transformed to align with the target LLM’s representational space. The resulting aligned hidden states $\mathbf{H}_{align}$ maintain dimensionality compatibility with the target LLM’s hidden states while extending the sequence length by one ($|\mathbf{X}_{input}|+1$): $\mathbf{H}_{align}=[\mathbf{FF}(\mathbf{H}_{enc});\mathbf{t}]$, where $\mathbf{FF}$ is the feed-forward network to align the multilingual encoder’s hidden states with dimension $\mathbf{d}_{e}$ to the target LLM’s hidden states with dimension $\mathbf{d}_{t}$, and $\mathbf{t}\in\mathbb{R}^{d_{t}}$ is the trainable token. Moreover, we employ a masking mechanism to mask any original padding tokens in $\mathbf{H}_{enc}$ to prevent their influence on the target LLM’s processing, ensuring the model focuses on meaningful tokens.
3.2 Training Pipeline
@@
Figure 2: Overview of tasks and datasets in our benchmark. Crosslingual datasets are marked with a blue shade.
@@
Stage 1: Alignment Training. The initial training stage aligns the multilingual encoder’s representations with the target LLM’s embedding space. Specifically, we optimize the connector parameters $\theta_{c}$ and the multilingual encoder parameters $\theta_{e}$ while keeping the target LLM’s parameters fixed, ensuring stable convergence. The training employs two complementary objectives: (1) A masked reconstruction task where we randomly mask $k\%$ of input tokens such that $\mathbf{X}_{input}=\texttt{mask}(\mathbf{X},k)$, training the model to recover the original sequence $\mathbf{X}_{lm}=\mathbf{X}$. (2) An autoregressive completion task that focuses on next-token prediction, where the model learns to generate the target sequence $\mathbf{X}_{lm}$ conditioned on the input context $\mathbf{X}_{input}$. The training objective for both tasks is formulated as language modeling objective to generate the target sequence $\mathbf{X}_{lm}$ given the input sequence $\mathbf{X}_{input}$. This objective enables local token-level alignment through masked reconstruction task where the model learns to predict the masked tokens by leveraging the context. In addition, it exploits global semantic alignment through autoregressive completion task that encourages the model to capture semantic information of the input sequences to generate the target sequence. As such, our training strategy learns to align the multilingual encoder’s representations with the target LLM’s embedding space while preserving important semantic information of multilingual input sequences. Our training process is conducted using the standard cross-entropy loss function.
@@
Stage 2: Representation Finetuning. The second stage improves text representations through a contrastive learning process, effectively teaching the model to distinguish between positive and negative examples. Our approach leverages both in-batch negatives sampled from the current training batch and hard-negative examples specifically curated to enhance model training. Additionally, we incorporate bidirectional attention mechanisms within the target LLM, following recent advances in LLM’s representation learning Muennighoff et al. (2024); BehnamGhader et al. (2024); Lee et al. (2024); Man et al. (2024). This bidirectional context modeling significantly enhances the quality of learned representations by enabling the model to capture both forward and backward dependencies in the input sequence. During this stage, we finetune all components of LUSIFER, including the target LLM, the multilingual encoder, and the connector parameters, to optimize the model’s representation quality for embedding-specific tasks. The goal of this stage is to improve the quality of text representations by leveraging the advanced embedding capabilities of the target LLM while maintaining the multilingual understanding provided by the multilingual encoder.
4 Experiment

### v2

@@ Abstract.
Recent advancements in large language models (LLMs) based embedding models have established new state-of-the-art benchmarks for text embedding tasks, particularly in dense vector-based retrieval. However, these models predominantly focus on English, leaving multilingual embedding capabilities largely unexplored. To address this limitation, we present LUSIFER, a novel zero-shot approach that adapts LLM-based embedding models for multilingual tasks without requiring multilingual supervision. LUSIFER’s architecture combines a multilingual encoder, serving as a language-universal learner, with an LLM-based embedding model optimized for embedding-specific tasks. These components are seamlessly integrated through a minimal set of trainable parameters that act as a connector, effectively transferring the multilingual encoder’s language understanding capabilities to the specialized embedding model. Additionally, to comprehensively evaluate multilingual embedding performance, we introduce a new benchmark encompassing 5 primary embedding tasks, 123 diverse datasets, and coverage across 14 languages. Extensive experimental results demonstrate that LUSIFER significantly enhances the multilingual performance across various embedding tasks, particularly for medium and low-resource languages, without requiring explicit multilingual training data. The code and dataset for training are available at: https://github.com/hieum98/lusifer
Large Language Models, Representation Learning, Multilingual Embeddings
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; July 13-18, 2025; Padua, Italyjournalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730029isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Language models
1. Introduction
@@
Text embeddings, which provide dense vector representations of textual content (Mikolov et al., 2013; Devlin et al., 2019), have become fundamental building blocks in modern natural language processing. These embeddings encode semantic information and serve as an important component for numerous downstream applications, ranging from information retrieval and document reranking to classification, clustering, and semantic textual similarity assessment. Recently, the significance of high-quality embeddings has been further amplified by their crucial role in retrieval-augmented generation (RAG) systems (Lewis et al., 2020b). RAG architectures enable large language models (LLMs) to dynamically access and integrate external or proprietary knowledge without the need for model parameter updates, substantially enhancing their adaptability and accuracy (Wang et al., 2023; Liu et al., 2024b; Gao et al., 2024).
@@
The evolution of embedding models has witnessed remarkable advancements, progressing from static word embeddings (Robertson et al., 2009) through contextualized representations (Reimers and Gurevych, 2019; Gao et al., 2021a; Ni et al., 2021a) to state-of-the-art LLM-based embedding models (Wang et al., 2024b) that harness the sophisticated semantic understanding capabilities of large language models. These developments have substantially enhanced performance across various embedding tasks (Luo et al., 2024), achieving unprecedented accuracy in semantic similarity and retrieval applications. However, a critical limitation remains: the predominant focus on English in LLM-based embedding models has created a significant disparity in multilingual capabilities. This gap is especially pronounced in medium and low-resource languages, where English-centric models exhibit substantial performance degradation due to insufficient language-specific training data (Wang et al., 2020; Thakur et al., 2024). While recent advances in multilingual embedding models, particularly those leveraging multilingual pre-trained architectures, have demonstrated promising results in multilingual embedding tasks (Li et al., 2023; Wang et al., 2024c; Chen et al., 2024), their reliance on explicit multilingual supervision for embeddings constrains their applicability primarily to languages with abundant training resources, leaving the challenge of true language-agnostic representation largely unaddressed.
@@
To address this challenge, we present LUSIFER, a novel zero-shot approach that adapts English LLM-based embedding models for multilingual tasks without requiring explicit multilingual supervision. Drawing inspiration from recent advances in multimodal integration (Liu et al., 2024a; Lu et al., 2024), LUSIFER employs a unique architecture that bridges the gap between multilingual understanding and specialized embedding capabilities. At its core, LUSIFER leverages the robust multilingual representations from XLM-R (Conneau et al., 2020) and introduces a learnable connector mechanism to interface with English-optimized LLM embedding models. This approach enables LUSIFER to effectively transfer the multilingual understanding of XLM-R to the target LLM while inheriting advanced embedding capabilities of the LLM. In this way, LUSIFER can achieve effective multilingual representation capabilities without requiring explicit multilingual training data. We conduct comprehensive evaluations of LUSIFER through extensive experiments across 123 diverse datasets spanning 14 languages, focusing on five fundamental embedding tasks: Classification, Clustering, Reranking, Retrieval, and Semantic Textual Similarity (STS). Our experimental results demonstrate that LUSIFER substantially enhances the performance of English-centric LLM-based embedding models, achieving average improvements of 3.19 points across all tasks, with particularly significant gains observed for medium and low-resource languages (up to 22.15 improvement). To validate LUSIFER’s broader applicability and cross-lingual capabilities, we extend our evaluation to cross-lingual tasks using four comprehensive datasets that encompass over 100 languages, including several critically low-resource languages. LUSIFER significantly outperforms existing English-centric embedding models by 5.75 on average in cross-lingual scenarios. These results demonstrate the effectiveness of our approach in enhancing multilingual representation capabilities without explicit multilingual supervision.
@@
The theoretical foundation for LUSIFER’s effectiveness lies in its ability to create a language-agnostic universal space through the integration of a multilingual encoder (Pires et al., 2019; Libovický et al., 2020). We hypothesize that this universal space serves as a bridge between different languages, enabling the target language model to process semantic information independently of the input language. By mapping these language-neutral representations to the target model’s input space, we conjecture that the target LLM can grasp the semantics of these representations, thereby improving the quality of output embeddings across multiple languages. This mechanism allows the model to become less dependent on the specific language of the input, enabling it to better capture semantic information for embedding tasks in languages it rarely encountered during pretraining. Our empirical analysis using t-SNE visualization supports this hypothesis.
@@
Figure 1. Overview of LUSIFER. Left: Align a multilingual encoder with the target English-centric LLM only using English data and a minimal set of trainable parameter. Center: End-to-end representation finetune through contrastive learning on English text-embedding tasks using LoRA. Right: During inference, LUSIFER successfully processes text-embedding tasks across multiple languages.
2. Related Work
2.1. English-centric Embedding Models
Text embedding models have experienced significant advancement in recent years, driven by the evolution of pre-trained language models. Early successes with BERT-based architectures, as demonstrated in Sentence-BERT (Reimers and Gurevych, 2019), SimCSE (Gao et al., 2021a), and DPR (Karpukhin et al., 2020), established the foundation for modern embedding approaches. The field has since progressed to leverage LLMs, with recent works (Wang et al., 2024b; Muennighoff et al., 2024; BehnamGhader et al., 2024; Lee et al., 2024; Man et al., 2024) demonstrating substantial improvements in embedding quality and task performance through the enhanced representational capacity of LLMs (Luo et al., 2024). However, these advances primarily benefit high-resource language applications, as most state-of-the-art LLM-based embedding models are derived from English-centric foundation models (Jiang et al., 2023; Touvron et al., 2023) and trained predominantly on English or high-resource language datasets (Wang et al., 2024b). This bias has resulted in a significant performance gap between high-resource and low-resource languages, limiting the global applicability of these models. Our proposed method, LUSIFER, addresses this limitation by enabling effective multilingual representation without multilingual training data.
2.2. Zero-shot Multilingual Embedding
Multilingual Embedding has evolved through several distinct methodological approaches, each addressing the fundamental challenge of bridging language gaps in embedding tasks. Early successful approaches relied on translation models to enable multilingual understanding (Liu et al., 2020; Shi et al., 2021; Zhang and Misra, 2022). While effective, these methods introduced operational complexity by requiring external translation systems, limiting their practical deployment and scalability. The emergence of multilingual pre-trained language models, particularly XLM-R (Conneau et al., 2020), opened new possibilities for multilingual transfer. Recent works have demonstrated promising results by fine-tuning such models with contrastive learning objectives on multilingual data (Wang et al., 2024c; Chen et al., 2024; Sturua et al., 2024). However, these approaches face two key limitations: they require substantial multilingual training data, and moreover, they do not exploit the sophisticated semantic representations afforded by contemporary English-centric LLM architectures, which have demonstrated superior performance in capturing nuanced semantic relationships.
Recent advances in aligning multilingual and English-centric representations could offer a solution. By combining independently pre-trained representations, a paradigm that has shown remarkable success in multimodal alignment research (Alayrac et al., 2022; Liu et al., 2024a; Lu et al., 2024), these works bridge the gap between visual encoders and language models to enhance visual comprehension. As such, similar principles can be applied to align multilingual representations with LLM-based semantic spaces. While related efforts have explored aligning multiple LLMs for improved reasoning capabilities in multilingual settings (Bansal et al., 2024; Yoon et al., 2024), these approaches primarily target generation tasks and typically require large-scale alignment data. Our work extends these efforts by focusing on embedding tasks and leveraging a minimal set of parameters to align multilingual and English-centric representations, enabling enhanced multilingual representation capabilities without requirement for large-scale multilingual training data.
2.3. Multilingual Embedding Benchmarks
The evaluation landscape for multilingual embedding models has historically been fragmented across various benchmarks, each with significant limitations. While existing benchmarks have made valuable contributions, they often exhibit constrained scope: MINERS (Winata et al., 2024) provides evaluation across multiple languages but is limited to classification and STS tasks with only 11 datasets; XNLI (Conneau et al., 2018), XQuAD (Artetxe et al., 2020), and SIB-200 (Adelani et al., 2024) offer broad language coverage but focus exclusively on classification tasks; and MTEB (Muennighoff et al., 2023), despite its diverse task selection, primarily addresses high-resource languages. To address these limitations, we introduce a comprehensive evaluation framework that encompasses 5 fundamental embedding tasks—Classification, Clustering, Reranking, Retrieval, and STS—across an extensive collection of 123 datasets spanning 14 languages. This holistic approach enables systematic evaluation across both task and language dimensions, providing unprecedented insights into models’ multilingual capabilities. Furthermore, our benchmark extends beyond traditional multilingual evaluation by incorporating cross-lingual tasks, featuring coverage of over 100 languages, including critically low-resource languages that have been historically underrepresented in existing benchmarks. This extensive coverage allows for a more nuanced understanding of embedding models’ performance across the global linguistic landscape.
3. Methodology
@@
Previous works demonstrate that representations of multilingual encoder models exhibit inherent language-agnostic properties, facilitating zero-shot multilingual transfer (Pires et al., 2019; Libovický et al., 2020). Building upon this foundation, we propose LUSIFER, an embedding framework that aligns a multilingual encoder model with a target English-centric LLM’s representational space, enabling the target to encode semantics across multiple languages without extensive multilingual training. This section details our architectural design and two-stage training process for LUSIFER.
3.1. Model Architecture
@@
Following successful approaches in multimodal alignment (Alayrac et al., 2022; Liu et al., 2024a; Lu et al., 2024), we implement the connector as a 2-layers feed-forward network, $\mathbf{FF}$, augmented with a single trainable token appended to the multilingual encoder’s hidden states. Formally, given input tokens $\mathbf{X}_{input}$ (with necessary padding), the multilingual encoder’s hidden states $\mathbf{H}_{enc}$ are transformed to align with the target LLM’s representational space. The resulting aligned hidden states $\mathbf{H}_{align}$ maintain dimensionality compatibility with the target LLM’s hidden states while extending the sequence length by one ($|\mathbf{X}_{input}|+1$): $\mathbf{H}_{align}=[\mathbf{FF}(\mathbf{H}_{enc});\mathbf{t}]$, where $\mathbf{FF}$ is the feed-forward network to align the multilingual encoder’s hidden states with dimension $\mathbf{d}_{e}$ to the target LLM’s hidden states with dimension $\mathbf{d}_{t}$, and $\mathbf{t}\in\mathbb{R}^{d_{t}}$ is the trainable token. Moreover, we employ a masking mechanism to mask any original padding tokens in $\mathbf{H}_{enc}$ to prevent their influence on the target LLM’s processing, ensuring the model focuses on meaningful tokens.
3.2. Training Pipeline
@@
Figure 2. Overview of tasks and datasets in our benchmark. Crosslingual datasets are marked with a blue shade.
@@
Stage 1: Alignment Training. The initial training stage aligns the multilingual encoder’s representations with the target LLM’s embedding space. Specifically, we optimize the connector parameters $\theta_{c}$ and the multilingual encoder parameters $\theta_{e}$ while keeping the target LLM’s parameters fixed, ensuring stable convergence. The training employs two complementary objectives: (1) A masked reconstruction task where we randomly mask $k\%$ of input tokens such that $\mathbf{X}_{input}=\texttt{mask}(\mathbf{X},k)$, training the model to recover the original sequence $\mathbf{X}_{lm}=\mathbf{X}$. (2) An autoregressive completion task that focuses on next-token prediction, where the model learns to generate the target sequence $\mathbf{X}_{lm}$ conditioned on the input context $\mathbf{X}_{input}$. The training objective for both tasks is formulated as language modeling objective to generate the target sequence $\mathbf{X}_{lm}$ given the input sequence $\mathbf{X}_{input}$. This objective enables local token-level alignment through masked reconstruction task where the model learns to predict the masked tokens by leveraging the context. In addition, it exploits global semantic alignment through autoregressive completion task that encourages the model to capture semantic information of the input sequences to generate the target sequence. As such, our training strategy learns to align the multilingual encoder’s representations with the target LLM’s embedding space while preserving important semantic information of multilingual input sequences. Our training process is conducted using the standard cross-entropy loss function. This stage aims to establish a strong alignment between the multilingual encoder and the target LLM, enabling the target LLM to effectively process multilingual representations produced by the multilingual encoder.
@@
Stage 2: Representation Finetuning. The second stage improves text representations through a contrastive learning process, effectively teaching the model to distinguish between positive and negative examples. Our approach leverages both in-batch negatives sampled from the current training batch and hard-negative examples specifically curated to enhance model training. Additionally, we incorporate bidirectional attention mechanisms within the target LLM, following recent advances in LLM’s representation learning (Muennighoff et al., 2024; BehnamGhader et al., 2024; Lee et al., 2024; Man et al., 2024). This bidirectional context modeling significantly enhances the quality of learned representations by enabling the model to capture both forward and backward dependencies in the input sequence. During this stage, we finetune all components of LUSIFER, including the target LLM, the multilingual encoder, and the connector parameters, to optimize the model’s representation quality for embedding-specific tasks. The goal of this stage is to improve the quality of text representations by leveraging the advanced embedding capabilities of the target LLM while maintaining the multilingual understanding provided by the multilingual encoder.
The two-stage training process enables LUSIFER to effectively align multilingual representations with the target LLM’s embedding space, enhancing the target LLM’s multilingual representation capabilities without requiring explicit multilingual supervision.
4. Experiment